What number is used to complete the square for the equation y^2 + 16 = -y?

a) -64
b) 64
c) .25
d) .-25

Answer by ankor@dixie-net.com(22740)

What number is used to complete the square for the equation y^2 + 16 = -y?
Arrange as quadratic equation
y^2 - y + 16 = 0
y^2 - y + ___ = -16
the third term is calculated, take the coefficient y and square it: .5^2 = .25
y^2 - y + .25 = -16 + .25
:
.25 completes the square